4.2 REQUIRED

Study the information given below and answer the following questions:

4.2.1 Calculate the project’s Internal Rate of Return (answer expressed to two

decimal places).

(6)

4.2.2 Should the company consider investing in the project? Why? (1)

INFORMATION

Demat Limited is considering an investment in a project that costs R260 000 and should result in

cash savings of R81 000 per year for five years. The company’s cost of capital is 15%.


1
Expert's answer
2021-08-25T17:03:45-0400

Solution:

4.2.1). IRR calculation:

The formula for IRR is as follows:

"∑\frac{Cashflow}{(1 + IRR)^{t} } - Initial \;investment = 0"


"\frac{81,000}{1 + IRR^{1} } + \frac{81,000}{1 + IRR^{2} } + \frac{81,000}{1 + IRR^{3} } + \frac{81,000}{1 + IRR^{4} } + \frac{81,000}{1 + IRR^{5} } - 260,000 = 0"


Use trial and error till you find the rate that will make NPV = 0


Using excel, IRR = "16.86 \%"


IRR = "16.86 \%"

 

4.2.2). Demat Limited should invest in the project since the IRR value is greater than the company’s cost of capital, which means the project will bring higher returns.